itne and U.S. Customary Systems Related on page i. Which is greater: 1 inch or 1 centimeter? ii. Which is greater: 1 pound or 1 kilogram? iii. Which is greater: 1 quartz or 1 liter? Read the section on How Are Units of Length, Area, Volume, and Capacity Related on 502. 1. ii. 1 cubic centimeter is the is equal tol cubic centimeter of a cube whose has a side has of ii. 1 cubic meter means v. True or False: A 4-cm-by-4-cm-by-4-cm cube has an area as 64 cm v. True or False: A square of 64 cm x 64 cm has an area as 64 cm i, True or False: Asquare of 8 cm × 8 cm has an area as 64 cm2 i. True or False: 64 square cm is equal to 8 cm. times times

answer the 10 questions please!
Answer #1

i) 1 inch ( 1inch = 2.54 cm )

ii) 1 kg ( 1 kg = 2.2046 pounds )

iii) 1 liter ( 1 quartz = 0.946 liter )

i) 1 cm3

ii) Volume, 1 cm

iii) 1m , 1m, 1m

iv) False, it has area = 6A2 = 6*4*4 = 96cm2

v) False, it has area = A2 = 64*64 = 4096cm2

vi) True

vii) False, 64 square cm = 64 cm2